OpenStudy (anonymous):

Would someone like to help me with intercepts for a medal? I'll post the problem below...

OpenStudy (anonymous):

\[x^2y-x^2+4y=0\] I need to find the x and y intercepts of this problem, but I'm not sure how to. I've tried replacing each variable with 0 to solve for the other variable, but I keep getting strange answers...

OpenStudy (dessyj1):

Can you show me what you got when you tried to find the x-intercept?

OpenStudy (anonymous):

I just got 0

OpenStudy (anonymous):

\[x^2y−x^2+4y=0\] I factored the equation... \[x^2(y-1)+4y=0\] then I replaced y with 0... \[x^2((0)-1)+4(0)=0\] solved it... \[x^2=0\] and squared... \[x=0\]

OpenStudy (dessyj1):

Why do you think that is weird? Because that is the answer.

OpenStudy (dessyj1):

Also, you cannot factor that equation

OpenStudy (dessyj1):

actually, you just factored the first two terms so that's okay.

OpenStudy (anonymous):

I just wasn't sure about the answer since I originally got a different answer and didn't know which was correct. Thanks for clarifying though. :)

OpenStudy (anonymous):

By the way, you wouldn't happen to know anything about the symmetry of a function would you?

OpenStudy (dessyj1):

When y=0, x=0. It is a point on the origin. (0,0)

OpenStudy (dessyj1):

So, the x-int is 0 and a y-int is 0

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I graphed the function earlier and that's what it showed as well

OpenStudy (dessyj1):

I do not know anything about the symmetry of a function. That is another question you will have to ask.